OpenStudy (kittiwitti1):

hyperbolic trigonometric functions : "If sinhx= (15/8) and x>0, find the other hyperbolic functions at x" CHECK ONLY...

OpenStudy (kittiwitti1):

\[\sinh{x}=\frac{15}{8},x\lt0\]\[\sinh{x}=\frac{e^{x}-e^{-x}}{2}=\frac{15}{8}...\]

zepdrix (zepdrix):

check only? Do you have some work to show or something? :

OpenStudy (kittiwitti1):

am I going in the right direction?

zepdrix (zepdrix):

Mmm.. yes this seems right. Solve for x! :)

OpenStudy (kittiwitti1):

\[\frac{15}{8}=\frac{e^{x}-e^{-x}}{2}\rightarrow2(15)=8(e^{x}-e^{-x})\]does this work?

zepdrix (zepdrix):

Sure, that's a good second step. Let's divide a 2 out of each side,\[\large\rm 4e^x-4e^{-x}=15\]

zepdrix (zepdrix):

The next step is a little tricky. I won't spill the beans if you want to figure it out though :) Any ideas? :)

OpenStudy (kittiwitti1):

\(4(e^{x}-e^{-x})=15\)?

zepdrix (zepdrix):

Nooo silly :P I distributed the 4, we don't want to move backwards hehe. We have to apply some algebra trick to get rid of this negative exponent.

OpenStudy (kittiwitti1):

\[4e^{x}-\frac{4}{e^{x}}=15\]

zepdrix (zepdrix):

So you rewrote it in another way, but he's still there... Hmm

OpenStudy (kittiwitti1):

Any way to rewrite the \(\large{e^{-x}}\) is what you mean?

zepdrix (zepdrix):

Well I like how you wrote it in the denominator. To get rid of denominator stuff, we multiply through by the value down there, ya? :)

zepdrix (zepdrix):

So in this case, we'll multiply both sides of our equation by e^x.

OpenStudy (kittiwitti1):

Oh, I see what you are doing... removing any fractional forms for easier solving.

zepdrix (zepdrix):

\[\large\rm e^x\left(4e^{x}-\frac{4}{e^{x}}\right)=15e^x\]

zepdrix (zepdrix):

Ya it'll get rid of that fraction business which is great.

OpenStudy (kittiwitti1):

\[4e^{x}\times e^{x}=4e^{2x}\]yes?

OpenStudy (kittiwitti1):

and then treat it like a quadratic with \(x\) value as \(e^{x}\)

zepdrix (zepdrix):

Yes, but it might actually be useful to keep the square on the outside. \[\large\rm 4e^x\cdot e^x=4(e^x)^2\]If it's not immediately obvious why, that's ok.

zepdrix (zepdrix):

Ahh ya there ya go :P too fast for me hehe

OpenStudy (kittiwitti1):

no, I get what you mean... LOL sorry

zepdrix (zepdrix):

\[\large\rm 4(e^x)^2-4=15e^x\]\[\large\rm 4(e^x)^2-15e^x-4=0\]So ya some sort of quadratic business going on.

OpenStudy (kainui):

Nonono don't do it this way! When determining cosine (on the unit circle) when only given sine we would do the following: \[\sin x = \frac{b}{c}\] Use the pythagorean theorem to relate: \[a^2+b^2=c^2\] so we would end up with: \[\cos x = \frac{\sqrt{c^2-b^2}}{c}\] then from here you get tangent from: \[\tan x = \frac{\sin x}{\cos x}\] --- Similarly here we do the analogous thing, except instead of moving along a circle \(a^2+b^2=c^2\) you're moving along a hyperbola \(a^2-b^2=c^2\)! \[\sinh x = \frac{b}{c}\] Use the hyperbola to relate: \[a^2-b^2=c^2\] so we would end up with: \[\cosh x = \frac{\sqrt{c^2+b^2}}{c}\] then from here you get tangent from: \[\tanh x = \frac{\sinh x}{\cosh x}\] --- So you should check (using the exponentials, plug and chug 'em in) that you have: \[\cosh^2 x -\sinh^2 t = 1\] Since this is where this stands from.

OpenStudy (kainui):

No offense @zepdrix I like you haha, but doing it your way is like doing this: \[\sin x = \frac{15}{8} \]\[\frac{e^{ix}-e^{-ix}}{2i}= \frac{15}{8} \]\[\cdots\] in order to determine cosine, which feels wrong and is slower both for circles and the hyperbolic way too!

zepdrix (zepdrix):

Ahhh crap good call :P Literally all of the other hyperbolic identities slipped out of my brain when she mentioned the exponential one lol

zepdrix (zepdrix):

We did a whole crap load of work to solve for x. That doesn't really help us figure out cosh and tanh and those things.

OpenStudy (kainui):

I guess just for the fun of it, might as well just lay out some popular identities that are worth confirming at least once in everyone's life: \[e^x =\cosh x + \sinh x\]\[1=\cosh^2 x - \sinh^2 x\] Also, here's what I'm getting, your method should end up at the same answer once you solve the quadratic I think. \[\cosh(x) = \frac{\sqrt{c^2+b^2}}{c} = \frac{17}{8}\]\[\tanh(x) = \frac{15}{17}\]

zepdrix (zepdrix):

Hmm neato :d